# Functional differential equations driven by càdlàg rough paths

### v2

@@ 3.4. RDEs with constant delay
Maybe the most prominent example of rough functional differential equations are RDEs with constant delays, cf. e.g. [FR06, NNT08, BMCR14, CE20, BBR20]. In the present subsection we consider the delayed rough differential equation
@@
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the rough differential equation with constant delay (3.4). Moreover, consider $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and fixed controlled paths $(\tilde{Y}_{j},\tilde{Y}_{j}^{\prime})\in\mathcal{V}^{p}_{Z_{\cdot-r_{j}}}([0,T];\mathbb{R}^{k})$, $j=1,\ldots,\ell$, with corresponding solution $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$.
    Suppose that $\|Y,Y^{\prime}\|_{X,p},\|\tilde{Y},\tilde{Y}^{\prime}\|_{X,p}\leq K$, for some $K>0$, and $\|Y_{j}\|_{p}$, $\|\tilde{Y}_{j}\|_{p}$, $\|Y_{j},Y_{j}\|_{X,p}$, $\|\tilde{Y}_{j},\tilde{Y}_{j}^{\prime}\|_{X,p}\leq L$, for some $L>0$, $j=1,\ldots,\ell$. Then, we have the estimate
    $|Y_{0}-\tilde{Y}_{0}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,p}$
    $\quad\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p}+\sum_{j=1}^{\ell}|Y_{j;0}-\tilde{Y}_{j;0}|+\sum_{j=1}^{\ell}\|Y_{j},Y_{j}^{\prime};\tilde{Y}_{j},\tilde{Y}_{j}^{\prime}\|_{Z_{\cdot-r_{j}},p},$
@@
(ii) Local estimate on $[ir_{1},(i+1)r_{1}]$, $i=0,\ldots,N-1$. To apply the continuity result presented in Theorem 2.5 on the subintervals $[ir_{1},(i+1)r_{1}]$, we need to ensure that the functionals satisfy the required assumptions. This is given for $(F,F^{\prime}),(\tilde{F},\tilde{F}^{\prime})$ in Lemma 3.4, and for $(F-\tilde{F},F^{\prime}-\tilde{F}^{\prime})$ we refer to the proof of Corollary 3.3 (ii), and write $C_{F-\tilde{F},i}$ for the corresponding constant. By Theorem 2.5, it then holds the estimate
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[ir_{1},(i+1)r_{1}]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[ir_{1},(i+1)r_{1}]}$
$\quad\lesssim|Y_{ir_{1}}-\tilde{Y}_{ir_{1}}|+|F_{ir_{1}}(Y)-\tilde{F}_{ir_{1}}(\tilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{p,[ir_{1},(i+1)r_{1}]}+C_{F-\tilde{F},i}$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p}+\sum_{j=1}^{\ell}|Y_{j;0}-\tilde{Y}_{j;0}|+\|Y_{j},Y_{j}^{\prime};\tilde{Y}_{j},\tilde{Y}_{j}^{\prime}\|_{Z_{\cdot-r_{j}},p}$
$\qquad+\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[0,ir_{1}]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[0,ir_{1}]},$
@@ 3.5. RDEs with variable delay
Rough differential equations with variable delays represent a slight generalization of RDEs with constant delays. More precisely, let us consider the rough differential equation with variable delay
@@
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the RDE with variable delay (3.6). Moreover, consider $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and a fixed controlled path $(\tilde{Y}_{\eta},\tilde{Y}_{\eta}^{\prime})\in\mathcal{V}^{p}_{Z_{\cdot-\eta(\cdot)}}([0,T];\mathbb{R}^{k})$ with corresponding solution $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$.
    Suppose that $\|Y,Y^{\prime}\|_{X,p}$, $\|\tilde{Y},\tilde{Y}^{\prime}\|_{X,p}\leq K$, for some $K>0$, and $\|Y_{\eta},Y_{\eta}\|_{X,p}$, $\|\tilde{Y}_{\eta},\tilde{Y}_{\eta}^{\prime}\|_{X,p}\leq L$, for some $L>0$. Then, we have the estimate
    $|Y_{0}-\tilde{Y}_{0}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,p}$
    $\quad\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p}+|Y_{\eta;0}-\tilde{Y}_{\eta;0}|+\|Y_{\eta},Y_{\eta}^{\prime};\tilde{Y}_{\eta},\tilde{Y}_{\eta}^{\prime}\|_{Z_{\cdot-\eta(\cdot)},p},$
@@ 4. Application to stochastic differential equations with delay
One main application of rough path theory is a pathwise and robust approach to stochastic differential equations, see e.g. [FH20]. In this section we show how a càdlàg martingale and its delayed version can be lifted to a joint random rough path in the spirit of stochastic Itô integration. Consequently, this allows to apply the results on rough functional differential equations, provided in Section 2, to Itô stochastic differential equations (SDEs) with constant delays.
@@ 4.1. Delayed martingales as rough paths
in the spirit of stochastic Itô integration. Notice, although $Z$ is a martingale, the stochastic process $X$ is, in general, not a martingale anymore and, thus, we cannot directly apply stochastic Itô integration to define a rough path lift of $X$. Moreover, if $Z$ is a semimartingale, it seems not obvious under which conditions the associated process $X$ is again a semimartingale.
@@
We assume w.l.o.g. that $j_{1}=0$, and write $r=r_{j_{2}}$, and $M=Z^{i_{1}}$, $\tilde{M}=Z^{i_{2}}$. For $n\in\mathbb{N}$, we define
$H^{n}_{t}:=\sum_{k=0}^{N_{n}-1}\tilde{M}_{t_{k}^{n}-r,t_{k+1}^{n}-r}\mathbf{1}_{(t_{k}^{n},t_{k+1}^{n}]}(t),\quad t\in[0,T],$
@@
$\int_{0}^{t}H^{n}_{s}\,\mathrm{d}M_{s}=\sum_{k=0}^{N_{n}-1}\tilde{M}_{t_{k}^{n}\wedge t-r,t_{k+1}^{n}\wedge t-r}M_{t_{k}^{n}\wedge t,t_{k+1}^{n}\wedge t}.$
@@
where $H:=\Delta_{\cdot}\tilde{M}_{\cdot-r}$, and []delimited-[][$\cdot$][ ⋅ ] denotes the quadratic variation. Using the localizing sequence $\tau_{m}=T\wedge\inf\{t:|\tilde{M}_{t}|\geq m\}$, $m\in\mathbb{N}$, and replacing $H^{n}$ by $H^{n}_{\cdot\wedge\tau_{m}}$ and $H$ by $H_{\cdot\wedge\tau_{m}}$, we may assume that the integrand is uniformly bounded, so that we can apply the dominated convergence theorem. Since $H^{n}\to H$ converges pointwise as $n\to\infty$, this shows (4.1).
@@
$\int_{0}^{t}H_{s}\,\mathrm{d}M_{s}=\sum_{s\leq t}\Delta_{s}\tilde{M}_{\cdot-r}\Delta_{s}M,$
@@
$4[M,\tilde{M}_{\cdot-r}]=\sum_{s\leq\cdot}(\Delta_{s}M+\Delta_{s}\tilde{M}_{\cdot-r})^{2}-\sum_{s\leq\cdot}(\Delta_{s}M-\Delta_{s}\tilde{M}_{\cdot-r})^{2}$
has càdlàg sample paths of finite variation, as both terms on the right hand side are monotonically increasing, which implies that $[M,\tilde{M}_{\cdot-r}]$ has almost surely finite $\frac{p}{2}$-variation, and concludes the proof. ∎
@@
The stochastic integral $\int_{0}^{t}Z^{i_{1}}_{u-r_{j_{1}}-}\,\mathrm{d}Z^{i_{2}}_{u-r_{j_{2}}}$ can be defined using classical stochastic Itô integration if $r_{j_{1}}>r_{j_{2}}$. Indeed, the stochastic process $(Z^{i_{2}}_{t-r_{j_{2}}})_{t\in[0,T]}$ is a martingale and the stochastic process $(Z^{i_{1}}_{t-r_{j_{1}}})_{t\in[0,T]}$ is predictable, both with respect to the filtration $(\mathcal{F}_{t-r_{j_{2}}})_{t\in[0,T]}$ with $\mathcal{F}_{t}:=\{\Omega,\emptyset\}$ for $t<0$.
@@
4.2. SDEs with delays as random RDEs
Let us consider the SDE with constant delays
@@ 4.2. SDEs with delays as random RDEs
    Let $(V,V^{\prime})$ be an adapted stochastic process such that $(V(\omega),V^{\prime}(\omega))\in\mathcal{V}^{p}_{X(\omega)}$ for almost every $\omega\in\Omega$, and assume that $\sup_{t\in[0,T]}|V^{\prime}(\omega)|_{t}$ is uniformly bounded in $\omega$. Then, the rough integral exists and coincides $\mathbb{P}$-almost surely with the stochastic Itô integral, that is
@@
Step 2. Since $V^{\prime}$ is assumed to be bounded, we are left to show that for $i,j=1,\ldots,d$,
@@
in probability, which then implies $\mathbb{P}$-almost sure convergence, along some subsequence. From here on, we take the sequence of partitions $\mathcal{P}^{n}=\{0=t^{n}_{0}<t^{n}_{1}<\ldots<t^{n}_{N_{n}}=T\}$, $n\in\mathbb{N}$, of the interval $[0,T]$ with vanishing mesh size, so that $|\mathcal{P}^{n}|\to 0$ as $n\to\infty$. Moreover, let $i>j$.
@@
While we considered square-integrable martingales and the associated stochastic differential equations with constant delays in this section, the presented results can be generalized in a fairly straightforward manner to:
@@
    SDEs/RDEs with variable delays of the form (3.6), as long as $\eta$ is assumed to be bounded with $\eta(t)\geq\varepsilon$, $t\in[0,T]$, for some $\varepsilon>0$.

### v3

@@ 3.4. RDEs with constant delay
Maybe the most prominent example of rough functional differential equations are RDEs with constant delay, cf. e.g. [FR06, NNT08, BMCR14, CE20, BBR20]. In the present subsection we consider the delayed rough differential equation
@@
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the rough differential equation with constant delay (3.4). Moreover, consider $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and fixed controlled paths $(\widetilde{Y}_{j},\widetilde{Y}_{j}^{\prime})\in\mathcal{V}^{p}_{Z_{\cdot-r_{j}}}([0,T];\mathbb{R}^{k})$, $j=1,\ldots,\ell$, with corresponding solution $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$.
    Suppose that $\|Y,Y^{\prime}\|_{X,p},\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p}\leq K$, for some $K>0$, and that $\|Y_{j}\|_{p}$, $\|\widetilde{Y}_{j}\|_{p}$, $\|Y_{j},Y_{j}\|_{X,p}$, $\|\widetilde{Y}_{j},\widetilde{Y}_{j}^{\prime}\|_{X,p}\leq L$, for some $L>0$, $j=1,\ldots,\ell$. Then, we have the estimate
    $|Y_{0}-\widetilde{Y}_{0}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p}$
    $\quad\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p}+\sum_{j=1}^{\ell}|Y_{j;0}-\widetilde{Y}_{j;0}|+\sum_{j=1}^{\ell}\|Y_{j},Y_{j}^{\prime};\widetilde{Y}_{j},\widetilde{Y}_{j}^{\prime}\|_{Z_{\cdot-r_{j}},p},$
@@
(ii) Local estimate on $[ir_{1},(i+1)r_{1}]$, $i=0,\ldots,N-1$. To apply the continuity result presented in Theorem 2.5 on the subintervals $[ir_{1},(i+1)r_{1}]$, we need to ensure that the functionals satisfy the required assumptions. This is given for $(F,F^{\prime}),(\widetilde{F},\widetilde{F}^{\prime})$ in Lemma 3.4, and for $(F-\widetilde{F},F^{\prime}-\widetilde{F}^{\prime})$ we refer to the proof of Corollary 3.3 (ii), and write $C_{F-\widetilde{F},i}$ for the corresponding constant. By Theorem 2.5, it then holds the estimate
$\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[ir_{1},(i+1)r_{1}]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[ir_{1},(i+1)r_{1}]}$
$\quad\lesssim|Y_{ir_{1}}-\widetilde{Y}_{ir_{1}}|+|F_{ir_{1}}(Y)-\widetilde{F}_{ir_{1}}(\widetilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{p,[ir_{1},(i+1)r_{1}]}+C_{F-\widetilde{F},i}$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p}+\sum_{j=1}^{\ell}|Y_{j;0}-\widetilde{Y}_{j;0}|+\|Y_{j},Y_{j}^{\prime};\widetilde{Y}_{j},\widetilde{Y}_{j}^{\prime}\|_{Z_{\cdot-r_{j}},p}$
$\qquad+\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[0,ir_{1}]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[0,ir_{1}]},$
@@ 3.5. RDEs with variable delay
Rough differential equations with variable delay represent a slight generalization of RDEs with constant delay. More precisely, let us consider the rough differential equation with variable delay
@@
    Let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ be the unique solution to the RDE with variable delay (3.6). Moreover, consider $(\tilde{y},\tilde{y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and a fixed controlled path $(\widetilde{Y}_{\eta},\widetilde{Y}_{\eta}^{\prime})\in\mathcal{V}^{p}_{Z_{\cdot-\eta(\cdot)}}([0,T];\mathbb{R}^{k})$ with corresponding solution $(\widetilde{Y},\widetilde{Y}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$.
    Suppose that $\|Y,Y^{\prime}\|_{X,p}$, $\|\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p}\leq K$, for some $K>0$, and $\|Y_{\eta},Y_{\eta}\|_{X,p}$, $\|\widetilde{Y}_{\eta},\widetilde{Y}_{\eta}^{\prime}\|_{X,p}\leq L$, for some $L>0$. Then, we have the estimate
    $|Y_{0}-\widetilde{Y}_{0}|+\|Y,Y^{\prime};\widetilde{Y},\widetilde{Y}^{\prime}\|_{X,p}$
    $\quad\lesssim|y_{0}-\tilde{y}_{0}|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p}+|Y_{\eta;0}-\widetilde{Y}_{\eta;0}|+\|Y_{\eta},Y_{\eta}^{\prime};\widetilde{Y}_{\eta},\widetilde{Y}_{\eta}^{\prime}\|_{Z_{\cdot-\eta(\cdot)},p},$
@@ 4. Application to stochastic differential equations with delay
One main application of rough path theory is a pathwise and robust approach to stochastic differential equations, see e.g. [FH20]. In this section we show how a càdlàg martingale and its delayed version can be lifted to a joint random rough path in the spirit of stochastic Itô integration. Consequently, this allows to apply the results on rough functional differential equations, provided in Section 2, to Itô stochastic differential equations (SDEs) with constant delay.
@@ 4.1. Delayed martingales as rough paths
in the spirit of stochastic Itô integration. Recall, for a martingale $(S_{t})_{t\in[0,T]}$, stochastic Itô integration allows to define the integral $\int_{0}^{t}\varphi_{s}\,\mathrm{d}S_{s}$ if $(\varphi_{t})_{t\in[0,T]}$ is a stochastic process with left-continuous sample paths with right-limits which is adapted to the augmented filtration generated by $(S_{t})_{t\in[0,T]}$. For a comprehensive introduction to stochastic integration see, e.g., [Pro05]. In the following, when writing a stochastic integral, like $\int_{0}^{t}\varphi_{s}\,\mathrm{d}S_{s}$, we will always implicitly refer to the augmented filtration generated by $(S_{t})_{t\in[0,T]}$ if not explicitly stated otherwise.
To construct a random rough path above the stochastic process $X=(X_{t})_{t\in[0,T]}$, the main challenge is to establish the existence of the random integral $\int_{0}^{t}Z_{t-r_{j_{1}}}\,\mathrm{d}Z_{t-r_{j_{2}}}$ for $j_{1}<j_{2}$ since $(Z_{t-r_{j_{1}}})_{t\in[0,T]}$ is, in general, not adapted to the augmented filtration generated by $(Z_{t-r_{j_{2}}})_{t\in[0,T]}$.
@@
We assume w.l.o.g. that $j_{1}=0$, and write $r=r_{j_{2}}$, and $M=Z^{i_{1}}$, $\widetilde{M}=Z^{i_{2}}$. For $n\in\mathbb{N}$, we define
$H^{n}_{t}:=\sum_{k=0}^{N_{n}-1}\widetilde{M}_{t_{k}^{n}-r,t_{k+1}^{n}-r}\mathbf{1}_{(t_{k}^{n},t_{k+1}^{n}]}(t),\quad t\in[0,T],$
@@
$\int_{0}^{t}H^{n}_{s}\,\mathrm{d}M_{s}=\sum_{k=0}^{N_{n}-1}\widetilde{M}_{t_{k}^{n}\wedge t-r,t_{k+1}^{n}\wedge t-r}M_{t_{k}^{n}\wedge t,t_{k+1}^{n}\wedge t}.$
@@
where $H:=\Delta_{\cdot}\widetilde{M}_{\cdot-r}$, and []delimited-[][$\cdot$][ ⋅ ] denotes the quadratic variation. Using the localizing sequence $\tau_{m}=T\wedge\inf\{t:|\widetilde{M}_{t}|\geq m\}$, $m\in\mathbb{N}$, and replacing $H^{n}$ by $H^{n}_{\cdot\wedge\tau_{m}}$ and $H$ by $H_{\cdot\wedge\tau_{m}}$, we may assume that the integrand is uniformly bounded, so that we can apply the dominated convergence theorem. Since $H^{n}\to H$ converges pointwise as $n\to\infty$, this shows (4.1).
@@
$\int_{0}^{t}H_{s}\,\mathrm{d}M_{s}=\sum_{s\leq t}\Delta_{s}\widetilde{M}_{\cdot-r}\Delta_{s}M,$
@@
$4[M,\widetilde{M}_{\cdot-r}]=\sum_{s\leq\cdot}(\Delta_{s}M+\Delta_{s}\widetilde{M}_{\cdot-r})^{2}-\sum_{s\leq\cdot}(\Delta_{s}M-\Delta_{s}\widetilde{M}_{\cdot-r})^{2}$
has càdlàg sample paths of finite $1$-variation, as both terms on the right hand side are monotonically increasing, which implies that $[M,\widetilde{M}_{\cdot-r}]$ has $\mathbb{P}$-almost surely finite $\frac{p}{2}$-variation, and concludes the proof. ∎
@@
The stochastic integral $\int_{0}^{t}Z^{i_{1}}_{u-r_{j_{1}}-}\,\mathrm{d}Z^{i_{2}}_{u-r_{j_{2}}}$ can be defined using classical stochastic Itô integration if $j_{1}>j_{2}$. Indeed, the stochastic process $(Z^{i_{2}}_{t-r_{j_{2}}})_{t\in[0,T]}$ is a martingale and the stochastic process $(Z^{i_{1}}_{t-r_{j_{1}}})_{t\in[0,T]}$ is predictable, both with respect to the filtration $(\mathcal{F}_{t-r_{j_{2}}})_{t\in[0,T]}$ with $\mathcal{F}_{t}:=\{\Omega,\emptyset\}$ for $t<0$.
@@
4.2. SDEs with delay as random RDEs
Let us consider the SDE with constant delay
@@ 4.2. SDEs with delay as random RDEs
    Let $(V,V^{\prime})$ be an adapted stochastic process such that $(V(\omega),V^{\prime}(\omega))\in\mathcal{V}^{p}_{X(\omega)}$ for almost every $\omega\in\Omega$. Then, the rough integral exists and coincides $\mathbb{P}$-almost surely with the stochastic Itô integral, that is
@@
Step 2. We are left to show that
(4.4) $\lim_{|\mathcal{P}|\to 0}\sum_{(s,t)\in\mathcal{P}}V^{\prime}_{s}\mathbb{X}_{s,t}=0,$
$\mathbb{P}$-almost surely, along some subsequence. It suffices to show that for $i,j=1,\dots,d$,
@@
in probability, which then implies $\mathbb{P}$-almost sure convergence, along some subsequence: if $V^{\prime}$ is $\mathbb{P}$-almost surely piecewise constant, then (4.5) implies (4.4). Otherwise, for any $\varepsilon>0$, there exists a suitable piecewise constant approximation ${V^{\prime}}^{,\varepsilon}$ of $V^{\prime}$ such that
$\|V^{\prime}-{V^{\prime}}^{,\varepsilon}\|_{\infty}\leq\varepsilon,$
$\mathbb{P}$-almost surely, see [AKLP23, Proposition B.1]. By a standard interpolation argument (e.g. [FV10, Proposition 5.5]), it follows, for any $q>p$, that
$\|V^{\prime}-{V^{\prime}}^{,\varepsilon}\|_{q}\leq\|V^{\prime}-{V^{\prime}}^{,\varepsilon}\|_{p}^{\frac{p}{q}}\|V^{\prime}-{V^{\prime}}^{,\varepsilon}\|_{\infty}^{1-\frac{p}{q}}\leq C\varepsilon^{1-\frac{p}{q}},$
$\mathbb{P}$-almost surely, where the implicit multiplicative constant $C$ is a random variable which does depend only on $p$, $q$ and $\|V^{\prime}\|_{p}$. Using [You36, (5.1)], we obtain that
$\Big{|}\sum_{(s,t)\in\mathcal{P}}V^{\prime}_{s}\mathbb{X}_{s,t}-\sum_{(s,t)\in\mathcal{P}}{V^{\prime}}^{,\varepsilon}_{s}\mathbb{X}_{s,t}\Big{|}\leq\Big{(}1+\zeta\Big{(}\frac{1}{q}+\frac{2}{p}\Big{)}\Big{)}\|V^{\prime}-{V^{\prime}}^{,\varepsilon}\|_{q}\|\mathbb{X}\|_{\frac{p}{2}}\leq C\varepsilon^{1-\frac{p}{q}},$
$\mathbb{P}$-almost surely for any partition $\mathcal{P}$, where the implicit multiplicative constant $C$ is a random variable which does depend only on $p$, $q$, $\|V^{\prime}\|_{p}$ and $\|\mathbb{X}\|_{\frac{p}{2}}$. Consequently, if
$\lim_{|\mathcal{P}|\to 0}\sum_{(s,t)\in\mathcal{P}}{V^{\prime}}^{,\varepsilon}_{s}\mathbb{X}_{s,t}=0,$
holds $\mathbb{P}$-almost surely, then so does (4.4), and it suffices to show (4.5).
Step 3. From here on, for the proof of (4.5), we consider the sequence of partitions $\mathcal{P}^{n}=\{0=t^{n}_{0}<t^{n}_{1}<\ldots<t^{n}_{N_{n}}=T\}$, $n\in\mathbb{N}$, of the interval $[0,T]$ with vanishing mesh size, so that $|\mathcal{P}^{n}|\to 0$ as $n\to\infty$. Moreover, let $i\geq j$.
@@
While we considered square-integrable martingales and the associated stochastic differential equations with constant delay in this section, the presented results can be generalized in a fairly straightforward manner to:
@@
    SDEs/RDEs with variable delay of the form (3.6), as long as $\eta$ is assumed to be bounded with $\eta(t)\geq\varepsilon$, $t\in[0,T]$, for some $\varepsilon>0$.